OpenStudy (anonymous):

to make sweet tea, ted uses 3 parts sugar to 5 parts tea. if ted needs 32 ounces of sweet tea, how many ounces of tea does he use?

OpenStudy (ash2326):

In 8 parts of tea 3 parts is sugar and 5 parts is tea so In 1 part of tea, \(\frac{3}{8}\) parts is sugar and \(\frac{5}{8}\) part is tea he needs to make 32 ounces of tea so for 32 parts(or ounces) of tea he needs \(\frac{3}{8} \times 32\) parts of sugar and \(\frac{5}{8}\times 32\) part of tea so he need tea= \(\frac{5}{8}\times 32\)=20 ounces of tea

OpenStudy (ash2326):

@amber1027 did you understand?

OpenStudy (anonymous):

Oh I see 8/32=4*5=20